SITE 1096

On leaving Site 1095, we deployed the proton precession magnetometer before crossing over the site. A continuous total field magnetic and 3.5-kHz profile was obtained that extended over and just to the southeast of Site 1096. We averaged 8.1 kt during the 58 nmi transit to Site 1096.

Hole 1096A
Hole 1096A was spudded at 1130 hr on 25 February with the APC at a seafloor depth estimated from recovery to be 3152.0 m (3163.2 mbrf). APC coring advanced to refusal at a depth of 140.7 mbsf (Core 15H) with 84.2% recovery. The bit cleared the seafloor at 0145 hr on 26 February, and the JOIDES Resolution was offset 20 m west of Hole 1096A.

Hole 1096B
Before spudding Hole 1096B, a water-sampling temperature probe (WSTP) sample of bottom water was obtained. Hole 1096B was spudded with the APC at 0700 hr on 26 February at a seafloor depth of 3152.5 m (3163.7 mbrf). Piston coring advanced to 108.3 mbsf (Core 12H) without incident but on Core 13H encountered a change in the formation and only advanced 7.6 m. We switched to XCB for the next two cores (115.9 to 135.2 mbsf) but got virtually no recovery. The hole was then deepened with the APC to refusal at 166.9 mbsf. The interval from about 110 to 170 mbsf, besides causing some coring difficulties, marks a zone where physical properties change and where a lithologic boundary occurs. Adara temperature tool measurements were obtained at the mudline, 32.2 mbsf (Core 4H), 60.8 mbsf (Core 7H), and 89.3 mbsf (Core 10H). Orientation data were obtained for Cores 4H, 5H, and 6H.

Coring continued in Hole 1096B with the XCB, advancing steadily to 212.5 mbsf, where a Davis-Villinger Temperature Probe (DVTP) was deployed on the wireline to measure formation temperature. Thereafter, circulation was poor and recovery reduced. At 260.6 mbsf, after a second DVTP deployment, the drill string was tripped, and three of the four bit nozzles were found to be plugged with clay and glacial dropstones. The nozzles were cleared and the bit seal replaced. From DVTP data and other data collected during APC coring, an initial temperature gradient of about 80°C/km was estimated.

Hole 1096C
Hole 1096C was started 20 m west of Hole 1096B by first drilling to 114.0 mbsf. We then alternated between coring and drilling in the interval from 114.0 to 231.4 mbsf, taking two APC and four XCB cores to fill gaps in recovery from the first two holes.

Continuous XCB coring began at 255 mbsf and advanced to 588.6 mbsf (Core 41X) without incident and with generally good recovery (82.7%). While cutting Core 42X (588.6 to 598.1 mbsf), the hard formation cutting shoe failed downhole. Because of the need to ensure that sediments producing a prominent seismic reflector had been sampled, an additional core (43X) was obtained (598.1 to 607.7 mbsf), in which a small piece of ground-up cutting shoe was recovered.

In preparation for logging, a 20-bbl high-viscosity mud sweep was circulated and the pipe pulled back to 97.3 mbsf. The TC tool string was deployed at 0115 hr on 4 March but was retrieved without logging because of an electronic problem. The lithodensity, porosity, and natural gamma ray (IPLT) tool was then run into the pipe at 0600 hr and encountered an obstruction at 353 mbsf. The hole was logged up from this depth to the bit, with a short repeat section. During logging, additional tight spots were observed at 170 and 240 mbsf. In many parts of the hole, the diameter exceeded 18 in.

At noon the tool was recovered, and a wiper trip was conducted from 97 to 578 mbsf. The region from 578 to 607 mbsf was washed and reamed. After a 20-bbl mud flush was circulated, the bit was pulled back to 356 mbsf. The IPLT was deployed again and was able to advance from 356 to 556 mbsf, where an obstruction prevented further penetration. The hole was logged up from 556 mbsf to the bit.

At midnight on 4 March, the first tool string was recovered and the second tool (GHMT) deployed. The GHMT logged the hole up from 510 to 356 mbsf and was recovered and rigged down by 0930 hr on 5 March. The drill string was recovered, but the positioning beacon failed to release and was therefore not retrieved. The vessel departed for Site 1097 at 1530 hr on 5 March.

To 178 Site 1097

To 178 Table of Contents

ODP Publications

ODP Home